# MBBQ: A Dataset for Cross-Lingual Comparison of Stereotypes in Generative LLMs

### v2

@@ 2 Related work
Closest to our work, there exist the KoBBQ (Jin et al., 2023) and CBBQ (Huang & Xiong, 2023) datasets, both adaptations of the BBQ dataset into Korean and Chinese respectively. CBBQ was created by prompting GPT-4 to complete samples that were designed by humans, which as Jin et al. (2023) note is subject to GPT-4’s limitations, including its own biases. On the other hand, Jin et al. (2023) use cultural transfer techniques in translating the BBQ dataset to Korean and extending it to fit the South Korean cultural context. In contrast, we include three other languages, only consider stereotypes that apply to all of them, separate bias from task performance, and compare model biases across languages.
@@ 3.1 Dataset format
MBBQ is a translation of a carefully curated subset of the BBQ dataset, an English bias benchmark for question answering that consists of 58,492 samples across nine bias categories (Parrish et al., 2022). We specifically decided on the BBQ dataset because it measures implicit stereotypes through multiple choice questions, without requiring classifiers or more powerful LLMs to evaluate generated text, since those models, if available in multiple languages, may introduce their own social biases. Further, we believe that MBBQ could be a useful resource for investigating bias mitigation techniques like those applicable to the BBQ dataset (Ma et al., 2023; Gallegos et al., 2024; Kaneko et al., 2024) for non-English languages, as well as the cross-lingual debiasing effects of these techniques.
@@ 4.1 Models and prompts
Given that the BBQ dataset was originally created to benchmark question answering systems, which are generally language models with a fine-tuned multiple choice head (Sap et al., 2019; Rogers et al., 2020; Parrish et al., 2022), we need to adapt the task slightly to fit our generative LLMs. We follow Jin et al. (2023) and do so via prompting. In particular, we use 5 different prompts (available in Appendix C), also translated following the process described in Section 3.2, to create instructions out of the context, question, and answer options.
@@ 4.2 Evaluation
If no answer can be detected in the model’s response we consider this as neither a biased nor a counter-biased answer. To determine whether scores differ significantly across models and languages we utilize the Kruskal–Wallis $H$ test. [^6] A non-parametric equivalent of one-way ANOVA, used for testing whether samples (of equal or different size) originate from the same distribution (Kruskal & Wallis, 1952).
@@ 5.1 Ability to answer multiple choice questions
Figure 3: Average accuracy across languages on control-MBBQ for questions with disambiguated and ambiguous contexts. Chance accuracy is 33% in all cases.
@@ 5.2 Cross-lingual comparison of biases in MBBQ
We again break down the results by those obtained in disambiguated contexts and those in ambiguous contexts—both are displayed in Table  2. In disambiguated contexts, we notice that Aya and GPT3.5 are highly accurate, leaving very little room for bias. The other models show significant bias in at least one language, and significant differences in bias across languages. Since MBBQ only includes stereotypes that hold across all languages, this shows that models are inconsistent cross-lingually in the biases they exhibit. Out of the models that show significant biases, both Llama and Zephyr are most biased in Spanish. Mistral is the only model that is most biased in Turkish instead, even though nearly all models are least accurate in Turkish. Out of these models that show bias, the least accurate model, Llama, shows least bias in Turkish, and the other three models show least bias in English.
@@ 5.3 Bias per category
Model Aya GPT 3.5 Llama Mistral Zephyr
#T 94 89 67 56 58
Lang. E D S T E D S T E D S T E D S T E D S T
AccD 94.3 91.8 91.3 85.5 87.5 85.9 82.4 73.8 36.8 39.3 43.0 38.5 75.5 67.2 71.6 44.0 82.3 76.3 68.5 42.4
BiasD 0.0050 -0.0017 0.0052 -0.0004 -0.0011 0.0002 -0.0074 0.0010 0.0119* 0.0195* 0.0294* 0.0097 0.0054 0.0109 0.0106 0.0454* 0.0023 0.0366* 0.0384* 0.0264*
AccA 18.1 10.5 8.7 11.8 84.2 82.1 83.9 74.6 58.2 39.4 35.3 30.0 79.7 73.4 76.3 63.0 50.1 24.9 41.5 39.9
BiasA 0.0356* 0.0438* 0.1088* 0.0531* -0.0107 0.0035 0.0080 0.0167* 0.0259* 0.0262* 0.0329* 0.0245 0.0373* 0.0503* 0.0691* 0.0468* 0.0853* 0.1233* 0.1302* 0.0400*
@@ A.1 Selection and translation of templates
First, we exclude the race, religion, and nationality bias categories, since these pertain to biases that are highly different across languages and cultures, and many biases in these categories are specific to the US (Jin et al., 2023). The nationality category includes some countries in which our target languages are spoken, and the religion category includes the most prominent religions in those countries, so those stereotypes will likely differ across our target languages (Levy et al., 2023). Finally, following Jin et al. (2023), we exclude templates that refer to individuals using proper names, as names cannot be translated, nor be expected to have the same (gender) associations across languages. Aside from the US-centric stereotypes that are targeted by the different templates, BBQ also makes use of US-specific (brand) names and terms, such as ‘calling 911’ and ‘the TSA’. We replace these names and terms by more international equivalents, also to avoid that they are translated literally. After replacing these terms, we ask native speakers to evaluate the stereotypes targeted by the remaining templates.
@@ Appendix B Models
(Üstün et al., 2024) is a multilingual generative LLM with 13B parameters that was fine-tuned to follow instructions in 101 languages, over half of which are considered low-resource. Aya is based on the mT5 model, and was only instruction finetuned on fully open-source multilingual datasets: the xP3x Dataset, and extension of the xP3 dataset (Muennighoff et al., 2022), a collection from the Data Provenance Initiative (Longpre et al., 2023), Share-GPT Command, and the Aya Collection and Aya Dataset (Üstün et al., 2024) collected specifically for Aya.
@@
 [^15] https://openai.com/blog/introducing-chatgpt-and-whisper-apis
is a proprietary generative LLM by OpenAI. Little is known about its architecture and training data. We access GPT-3.5 Turbo through its API.
@@
(Tunstall et al., 2023) is a 7B parameter generative LLM that is a fine-tuned version of Mistral. In particular, it has been fine-tuned on two dialogue datasets, the UltraChat (Ding et al., 2023) and UltraFeedback (Cui et al., 2023) dataset which have been collected from open-source and proprietary LLMs. Zephyr benefits from distilled direct preference optimization, which utilizes the LLM outputs ranked by a teacher model from the UltraFeedback dataset. We use version $\beta$ of Zephyr 7B.

### v3

@@ 2 Related work
Closest to our work, there exist the KoBBQ (Jin et al., 2024) and CBBQ (Huang & Xiong, 2024) datasets, both adaptations of the BBQ dataset into Korean and Chinese respectively. CBBQ was created by prompting GPT-4 to complete samples that were designed by humans, which as Jin et al. (2024) note is subject to GPT-4’s limitations, including its own biases. On the other hand, Jin et al. (2024) use cultural transfer techniques in translating the BBQ dataset to Korean and extending it to fit the South Korean cultural context. In contrast, we include three other languages, only consider stereotypes that apply to all of them, separate bias from task performance, and compare model biases across languages.
@@ 3.1 Dataset format
MBBQ is a translation of a carefully curated subset of the BBQ dataset, an English bias benchmark for question answering that consists of 58,492 samples across nine bias categories (Parrish et al., 2022). We specifically decided on the BBQ dataset because it measures implicit stereotypes through multiple choice questions, without requiring classifiers or more powerful LLMs to evaluate generated text, since those models, if available in multiple languages, may introduce their own social biases. Further, we believe that MBBQ could be a useful resource for investigating bias mitigation techniques like those applicable to the BBQ dataset (Ma et al., 2024; Gallegos et al., 2024; Kaneko et al., 2024) for non-English languages, as well as the cross-lingual debiasing effects of these techniques.
@@ 4.1 Models and prompts
Given that the BBQ dataset was originally created to benchmark question answering systems, which are generally language models with a fine-tuned multiple choice head (Sap et al., 2019; Rogers et al., 2020; Parrish et al., 2022), we need to adapt the task slightly to fit our generative LLMs. We follow Jin et al. (2024) and do so via prompting. In particular, we use 5 different prompts (available in Appendix C), also translated following the process described in Section 3.2, to create instructions out of the context, question, and answer options. All reported results are averages across these 5 prompts.
@@ 4.2 Evaluation
If no answer can be detected in the model’s response we consider this as neither a biased nor a counter-biased answer. Those responses are discarded and not included in the average across prompts for that model’s bias score calculations. To determine whether scores differ significantly across models and languages we utilize the Kruskal–Wallis $H$ test. [^6] A non-parametric equivalent of one-way ANOVA, used for testing whether samples (of equal or different size) originate from the same distribution (Kruskal & Wallis, 1952).
@@ 5.1 Ability to answer multiple choice questions
Table 2: Average accuracy across languages on control-MBBQ for questions with disambiguated and ambiguous contexts. Chance accuracy is 33% in all cases.
@@ 5.2 Cross-lingual comparison of biases in MBBQ
We again break down the results by those obtained in disambiguated contexts and those in ambiguous contexts—both are displayed in Table  3. In disambiguated contexts, we notice that Aya and GPT3.5 are highly accurate, leaving very little room for bias. The other models show significant bias in at least one language, and significant differences in bias across languages. Since MBBQ only includes stereotypes that hold across all languages, this shows that models are inconsistent cross-lingually in the biases they exhibit. Out of the models that show significant biases, both Llama and Zephyr are most biased in Spanish. Mistral is the only model that is most biased in Turkish instead, even though nearly all models are least accurate in Turkish. Out of these models that show bias, the least accurate model, Llama, shows least bias in Turkish, and the other three models show least bias in English.
@@
Overall, we observe significant cross-lingual differences for the most biased and least accurate models. In disambiguated contexts, where the context contains enough information to answer the question, we can observe the difference between the models that rely on this information, and those that rely on their own biases instead. The former (Aya and GPT3.5) are highly accurate and unbiased, but the latter (Llama, Mistral, and Zephyr) are more likely to make biased choices whenever they are not accurate. They do so to a different extent depending on the language they were prompted in. In ambiguous contexts, we can differentiate between models admitting that there is not enough information in the context to answer the question, and those that are instead motivated by their own biases to choose an answer. Here, all models are biased in at least one language, whereas Aya, Mistral, and Zephyr are biased in all four. Therefore, in all languages, models are more likely to rely on their own biases when they cannot find an answer in the context. We also observe similarities across context types: Models are generally most biased in Spanish, the more accurate models are least biased in English, and the less accurate models are least biased in Turkish.
@@ 5.3 Bias per category
Model Aya GPT 3.5 Llama
#T 94 89 67
Lang. E D S T E D S T E D S T
AccD 94.3 91.8 91.3 85.5 87.5 85.9 82.4 73.8 36.8 39.3 43.0 38.5
BiasD 0.0050 -0.0017 0.0052 -0.0004 -0.0011 0.0002 -0.0074 0.0010 0.0119* 0.0195* 0.0294* 0.0097
AccA 18.1 10.5 8.7 11.8 84.2 82.1 83.9 74.6 58.2 39.4 35.3 30.0
BiasA 0.0356* 0.0438* 0.1088* 0.0531* -0.0107 0.0035 0.0080 0.0167* 0.0259* 0.0262* 0.0329* 0.0245
Model Mistral Zephyr
#T 56 58
Lang. E D S T E D S T
AccD 75.5 67.2 71.6 44.0 82.3 76.3 68.5 42.4
BiasD 0.0054 0.0109 0.0106 0.0454* 0.0023 0.0366* 0.0384* 0.0264*
AccA 79.7 73.4 76.3 63.0 50.1 24.9 41.5 39.9
BiasA 0.0373* 0.0503* 0.0691* 0.0468* 0.0853* 0.1233* 0.1302* 0.0400*
@@ A.1 Selection and translation of templates
First, we exclude the race, religion, and nationality bias categories, since these pertain to biases that are highly different across languages and cultures, and many biases in these categories are specific to the US (Jin et al., 2024). The nationality category includes some countries in which our target languages are spoken, and the religion category includes the most prominent religions in those countries, so those stereotypes will likely differ across our target languages (Levy et al., 2023). Finally, following Jin et al. (2024), we exclude templates that refer to individuals using proper names, as names cannot be translated, nor be expected to have the same (gender) associations across languages. Aside from the US-centric stereotypes that are targeted by the different templates, BBQ also makes use of US-specific (brand) names and terms, such as ‘calling 911’ and ‘the TSA’. We replace these names and terms by more international equivalents, also to avoid that they are translated literally. After replacing these terms, we ask native speakers to evaluate the stereotypes targeted by the remaining templates.
@@ Appendix B Models
(Üstün et al., 2024) is a multilingual generative LLM with 13B parameters that was fine-tuned to follow instructions in 101 languages, over half of which are considered low-resource. Aya is based on the mT5 model, and was only instruction finetuned on fully open-source multilingual datasets: the xP3x Dataset, and extension of the xP3 dataset (Muennighoff et al., 2023), a collection from the Data Provenance Initiative (Longpre et al., 2023), Share-GPT Command, and the Aya Collection and Aya Dataset (Üstün et al., 2024) collected specifically for Aya.
@@
is a proprietary generative LLM by OpenAI. [^15] https://openai.com/blog/introducing-chatgpt-and-whisper-apis Little is known about its architecture and training data. We access GPT-3.5 Turbo through its API. All responses from GPT-3.5 Turbo were collected between 3-1-2024 and 3-7-2024.
@@
(Tunstall et al., 2023) is a 7B parameter generative LLM that is a fine-tuned version of Mistral. In particular, it has been fine-tuned on two dialogue datasets, the UltraChat (Ding et al., 2023) and UltraFeedback (Cui et al., 2024) dataset which have been collected from open-source and proprietary LLMs. Zephyr benefits from distilled direct preference optimization, which utilizes the LLM outputs ranked by a teacher model from the UltraFeedback dataset. We use version $\beta$ of Zephyr 7B.